OpenStudy (anonymous):

Sam and garret worked over the summer. Together they earned a total of $425.00. Garret earned $25.00 more than Sam. How much did each of them earn? a. write a system of two equations with two variables to model this problem. b. Use substitution or the elimination method to solve the system.

OpenStudy (anonymous):

\(\bbox[9pt,#f9bec7,border:7pt solid #33cccc]{\color{white}{\Huge\cal\dagger Hello~:) \dagger}}\)

OpenStudy (anonymous):

S+G=(t+25) assuming s=sams amount and g=garrett's amount and t= total amount, s=200 and g=250